- Press Release
- August 15, 2022
NOAA GOES SATOPS Morning Report 30 Mar 2001
No change in the status of the POES spacecraft; instrument configuration
changes noted below.
NOAA-11 SBUV anomaly actions continue. The SBUV grating drive
has remained in an anomalous state since
approximately 0243Z on March 27. Commanding actions included:
1. Rev 64515 / W at 1601Z: Alternated between SBUV flex memory
2 and flex memory 0. There appears to be no
improvement in the SBUV.
2. Rev 64518 / F at 2059Z: SBUV testing continues. Flex
memory 0 Position mode was selected at the end of this pass,
leaving the grating parked at position 501.
POES operations were nominal over the past 24 hours.
No change in the status of the DMSP spacecraft.
DMSP operations were nominal over the past 24 hours.
No change in the status of the GOES spacecraft.
All eclipse operations nominal. No frame breaks noted.
To review the eclipse schedules for the GOES-8 and GOES-10 spacecraft,
http://www.ssd.noaa.gov/SSD/GOES/eclipse.html to identify images canceled
to satisfy eclipse or instrument keep-out-zone (KOZ) requirements.
Today is the last day spacecraft contractor, SS/L, is in town this week
for a series of COP reviews and launch sims.
GOES operations were nominal over the past 24 hours.
Questions or comments to Tim.Walsh@noaa.gov